OpenStudy (anonymous):

The height of an equilateral rectangular prism decreases by two units. The new lateral area is less than the original by how much? Answer two times the area of the base two times the perimeter of the base two less than the the area of the base two less than the perimeter of the base I think it is C but im not sure

OpenStudy (darthsid):

What's the formula for the area of the prism?

OpenStudy (anonymous):

mmm im not sure

OpenStudy (anonymous):

b*h

OpenStudy (darthsid):



OpenStudy (darthsid):

That's a rectangular prism. The lateral surface area is the area of the 4 sides, not including the base and the top.

OpenStudy (darthsid):

so the formula for the lateral surface area is Perimeter of base X height

OpenStudy (darthsid):

Now, if the new height is decreased by two, so it is h-2 so the area is : P x (h-2)

OpenStudy (darthsid):

are you understanding so far?

OpenStudy (anonymous):

yessss

OpenStudy (darthsid):

\[P \times ( h-2) = P \times h - P \times 2\]

OpenStudy (anonymous):

so it would be D ?

OpenStudy (darthsid):

But, P x h is the area of the original prism! the above equation basically says: New Area = Old Area - P x 2

OpenStudy (darthsid):

No, P x 2 means Perimeter times 2

OpenStudy (anonymous):

b?

OpenStudy (darthsid):

yeah, it's B

OpenStudy (anonymous):

thank you once again : )

OpenStudy (darthsid):

Glad to help, good luck!